Bill and Ben live 200km apart. They begin cycling towards each other at the same time, Bill traveling at 10km/h and Ben at 15km/h. Find the distance each will have travelled when they meet.
Answer by Alan3354(69443)

They travel 200 km at a combined speed of 25 km/hr (10 + 15)
200/25 = 8 hours pedalling.
8*10 = 80 km for Bill and
8*15 = 120 km